# M2D2: Exploring General-purpose Audio-Language Representations Beyond CLAP

### v1

@@ IV-A1 Basic setup
We followed M2D for the basic setup and implemented M2D2 on top of the M2D codebase as a variant of M2D-X[23]. Specifically, the audio encoder is ViT-base[39] with a 768-d feature, input audio duration of 6 s, and patch size of . The audio projector is a single-block transformer encoder with a 768-d feature, a single head, a 768-d feed-forward network latent feature, and a class token. It inputs audio features encoded by the audio encoder and outputs the class token feature as a summarized CLAP audio feature.
The audio samples were preprocessed to a log-scaled mel spectrogram with a sampling frequency of 16,000 Hz, window size of 25 ms, hop size of 10 ms, and 80 mel-spaced frequency bins in the range of 50 to 8000 Hz. We standardized spectrograms with the statistics of the first stage pre-training dataset (e.g., average of and standard deviation of ).
@@ IV-A2 Pre-training setup
In the additional stage 1.1, we fine-tuned the first-stage audio encoder in the same way as in the fine-tuning experiments in Section IV-B3. In the additional stage 2.1, we pre-trained the second-stage M2D2 with the same setting as in the second stage once again, except for a masking ratio of 0.0 (i.e., no masking) and a limited dataset.
@@ IV-A3 Pre-training datasets
In the first stage, the pre-training data consists of AudioSet[17], VGGSound[45], and WavCaps[7], 2,532,215 samples in total. We made it closer to CLAP[6]; they used 4.6 M samples from AudioSet, WavCaps, and many other datasets. We used Auto-ACD[46] for the caption data of AudioSet and VGGSound and supplemented the missing captions with ”The sound of labels” using their labels. Since the curated dataset included data samples for the downstream tasks (AudioCaps[41], Clotho[47], ESC-50[16], US8K[48]) used in the evaluations, we removed all of these data samples. For the audio data, we randomly cropped 6-s segments during the training.
@@
under a unified linear evaluation except Pengi, Cacophony, and CoLLAT.
@@
| (Previous studies: CLAP and audio-language models) | |||||||||||
@@
| (Proposed: Audio-language models) | |||||||||||
| M2D2, M2D2-, M2D2+ | 95.2 0.3 | 89.5 0.2 | 95.6 0.1 | 68.6 0.5 | 98.1 0.1 | 73.4 0.7 | 86.6 1.6 | 76.5 0.4 | 42.2 0.3 | 80.6 | |
| M2D2 | ✓ | 97.9 0.5 | 89.7 0.3 | 94.8 0.2 | 56.3 0.5 | 97.1 0.0 | 69.7 0.3 | 86.3 1.8 | 76.7 0.4 | 41.1 0.3 | 78.8 |
| (Ablation using a non-LLM text encoder and AudioSet only as audio data. The underlined results outperform M2D2-) | |||||||||||
| M2D20-(M2D-CLAP[11]) | 96.3 0.3 | 88.8 0.6 | 95.8 0.3 | 70.3 0.4 | 98.3 0.1 | 73.4 0.2 | 84.1 1.5 | 78.0 0.5 | 42.4 0.6 | 80.8 | |
@@ IV-A4 Evaluation tasks
IV-B Evaluation: Audio Features
We evaluated the effectiveness of the pre-trained audio features under the frozen feature and fine-tuning settings on general tasks and the frozen feature setting on music tasks. We followed [15][23] for the frozen feature (linear evaluation) and fine-tuning experimental setup using a unified evaluation platform, EVAR444https://github.com/nttcslab/eval-audio-repr, on general tasks, and used MARBLE benchmark[59] for the music task evaluation.
IV-B1 Frozen audio features on general audio tasks
Experimental setup
Results: Frozen audio features on general audio tasks
The results of the linear evaluation shown in Table II indicate that supervised learning and CLAP models perform poorly on some tasks, whereas the M2D2/M2D variants generally perform well. For example, supervised learning models AST, HTS-AT, and PANNs show 7.8 to 18.1 on VC1, and CLAP models such as LAION-CLAP show 8.4 to 27.2 on Surge, whereas the M2D variants show 52.7 to 73.4 on VC1 and 40.5 to 42.4 on Surge; we also observe these differences in other speech tasks. Note that the CLAP models use supervised learning models as their audio encoders; hence, the results indicate that applying audio features learned using labels/captions to tasks outside the coverage of these signals is challenging. For example, a training signal of ”Male speech, man speaking” does not cover who is speaking on VC1 (speaker identification).
BEATs and CED, pre-trained with distillation of the previous training cycle models, show a similar trend to supervised learning models. CED shows 35.2 on VC1, and BEATs shows 35.7 on Surge, for example. These models have distilled the pre-trained models on AudioSet, making them focused more on AudioSet data distribution. The results also indicate that their features become difficult to apply for tasks with different data distributions.
In contrast, M2D2/M2D variants and an SSL model ATST perform well in all tasks. M2D2 improves performance on ESC-50 at 94.9 and US8K at 89.1 compared to M2D (note that M2D2-/M2D2+ share the same audio encoder with M2D2). However, it exhibits slight performance degradation in the speech task, especially VC1, indicating a trade-off between M2D and CLAP’s training signal.
Notably, M2D2 (M2D2 encoder fine-tuned to AudioSet) does not degrade performance severely in speech tasks while improving performance on ESC-50 and US8K. It shows SOTA performance of 97.9 on ESC-50 and 89.7 on US8K. However, it degrades performance to 56.3 on VC1 from M2D2’s 68.6, while it is still in the good performance range compared to previous models. Similarly, M2D, a fine-tuned version of the original M2D, improves performance on ESC-50/US8K and degrades on others slightly more than M2D2 but not as severely as the supervised learning models. These results show that the M2D pre-training helps learn general-purpose representations that do not severely compromise performance against subsequent fine-tuning.
In addition, M2D2, which is pre-trained only on AudioSet using GTE-Base[60] (non-LLM-based sentence embedding model) as a text encoder, performs better than M2D2 in most tasks, indicating that the pre-training setting, especially relying solely on AudioSet is beneficial for performing well with frozen features. However, as shown in Section IV-C, its CLAP features severely underperform M2D2.
In summary, the experiments validated the effectiveness of M2D2 and the variants as a frozen audio extractor on general audio tasks.
@@ IV-B1 Frozen audio features on general audio tasks
| (Previous studies: Audio and audio-language models) | |||||||||||||||||
@@
| (Proposed: Audio-language models) | |||||||||||||||||
| M2D2, M2D2-, M2D2+ | 91.5 | 40.8 | 82.6 | 59.3 | 76.7 | 77.4 | 89.1 | 77.4 | 91.8 | 79.3 | 21.7 | 78.0 | 15.9 | 87.9 | 20.8 | 83.7 | 30.3 |
| M2D2 | 91.8 | 41.6 | 85.9 | 58.2 | 76.2 | 79.7 | 89.5 | 78.9 | 92.7 | 79.7 | 22.2 | 78.8 | 15.8 | 88.3 | 20.9 | 84.0 | 30.9 |
| (Ablation using a non-LLM text encoder and AudioSet only as audio data. The underlined results outperform M2D2-) | |||||||||||||||||
| M2D20-(M2D-CLAP[11]) | 91.6 | 41.0 | 84.3 | 61.9 | 77.4 | 80.6 | 88.8 | 78.1 | 92.3 | 79.0 | 21.6 | 78.2 | 16.0 | 88.0 | 20.8 | 83.8 | 30.4 |
@@ IV-B2 Frozen audio features on music tasks
To further assess the frozen feature, we evaluated models on the music audio tasks with various music audio signals.
@@
We used the MARBLE benchmark[59], which features diverse music audio tasks for music tracks, instrument sounds, and vocals. We used the tasks for which we could reproduce the results in the paper with the released code555https://github.com/a43992899/MARBLE-Benchmark, including music tagging and genre classification on MagnaTagATune (MTT)[64], GTZAN[53], and MTG-Jamendo (MTG)[65], emotion detection on Emomusic[66] and MTG, pitch and instrument (inst) classification on NSynth[54] and MTG, and vocal technique detection (tech) and singer identification on VocalSet[67].
The evaluation metrics are the area under the ROC curve (ROC) and the average precision (AP), except for Emomusic, which uses the determination coefficient between the model’s regression results and human annotations of valence (R2V) and arousal (R2A)[66]. The results are the averages of five tests with different random seeds.
@@
The results of the music task benchmark in Table III show that general audio models perform comparably to the SOTA music audio models. They even outperform music models on GTZAN and NSynth (Inst) while underperforming them in pitch classification NSynth (Pitch). While music models are pre-trained on large-scale music databases, general models are also pre-trained using large-scale music data because more than half (1M) of the AudioSet samples are tagged as ”music”[17]. In addition, most of the music and general models share a similar SSL-based training scheme. These facts would explain the comparable performance of general models on the music-specific benchmark. However, the CLAP model, CLAP2023, using a supervised learning audio backbone, underperforms other general and music models except for NSynth Inst666NSynth is contained in the CLAP2023 training data., suggesting the need for SSL to generalize in music tasks. In contrast, the SSL general audio models perform well in music tasks, indicating that pre-training on AudioSet with SOTA SSL methods can generalize well.
The models pre-trained using both SSL and supervised learning of AudioSet (M2D, M2D2, and BEATs) show stronger performance, and M2D2 demonstrates new SOTA results over the SOTA music audio models. In addition, as in the frozen audio feature evaluation on general tasks, M2D2 performs better than M2D2 in most music tasks, also indicating that pre-training solely on AudioSet is beneficial for performing well with frozen features. To summarize, the experiment confirmed that M2D2 models perform well and show SOTA performance in diverse music tasks.
@@ IV-B3 Fine-tuning
We followed M2D[23], ATST[32], and Audio-MAE[31] for the fine-tuning procedure. Specifically, a linear classifier was added on top of the pre-trained model to train the entire network. We used SpecAugment[68], Mixup[69, 15], and Random Resize Crop (RRC)[15] for data augmentation, along with Structured Patchout[70] that randomly masks patches during training. We used SGD, LARS, AdamW optimizers, and cosine annealing[71] learning rate scheduling. We used the hyperparameters for each task as listed in Table IV. For Structured Patchout, we used 768-d features calculated by averaging all ViT output features because Eq. (2) is not applicable to the masked patches. We used the tasks commonly used in previous studies, namely ESC-50, SPCV2, and VC1, the same as in the linear evaluation, plus AS2M and AS20K.
@@
| (Proposed: Audio-Language models) | |||||
| M2D2, M2D2-, M2D2+ | 49.0 0.1 | 42.1 0.0 | 97.8 0.1 | 98.4 0.0 | 95.5 0.3 |
| M2D2 | - | 46.7 0.2 | 98.5 0.0 | 98.4 0.1 | 94.0 0.2 |
| (Ablation using a non-LLM text encoder and AudioSet only as audio data) | |||||
| M2D20-(M2D-CLAP[11]) | 48.5 0.1 | 41.8 0.2 | 97.4 0.2 | 98.3 0.1 | 95.5 0.2 |
@@
The experimental results in Table V show that M2D2 variants further improve the performance of the baseline M2D. Unlike the linear evaluation that evaluated all models except a few, we evaluated only our models in fine-tuning because reproducing the best results of the previous methods is quite demanding.
M2D2 improves the performance from M2D, which shows SOTA results of 49.0 for AS2M and top-level results of 42.1 for AS20K and 97.8 for ESC-50. On the other hand, VC1 performance is 95.5, which is lower than the original M2D’s performance of 96.3, suggesting a trade-off for learning CLAP. M2D2 with an audio encoder fine-tuned to AudioSet exhibits an even stronger trend, showing SOTA results of 46.7 for AS20K and 98.5 for ESC-50 while 94.0 for VC1, a -1.5 pp degradation compared to M2D2.
The experiment confirmed that M2D2 achieves audio features that exhibit SOTA performance. Notably, it shows top performance with a single pre-training, unlike other SOTA models that require multi-iteration/model pre-training. In addition, the performance boost of M2D2 from M2D20- (M2D-CLAP) suggests the benefit of using a powerful LLM-based sentence encoder.
@@ IV-C1 Zero-shot Classification
We conducted experiments to evaluate the zero-shot classification performance of CLAP features. Under the zero-shot classification setting, a model infers the class label of audio by picking a class with the highest similarity between a CLAP audio feature and CLAP text feature of each class.
@@
We chose the tasks found in previous papers, which include ESC-50 (ESC), US8K, CREMA-D (CRD), GTZAN (GTZ), NSynth (NS), AudioSet (AS), and FSD50K (FSD). We used AudioSet and FSD50K as references because M2D2 was pre-trained on AudioSet and part of FSD50K in WavCaps. We conducted experiments using a standard procedure to infer a class label of audio with the closest cosine distance between the features of each test sample and the label’s caption, and we obtained accuracies from the inference results. Table VI summarizes the rules for converting task labels to the captions. We used the evaluation platform EVAR, as in Section IV-B, to evaluate M2D2 and pervious models.
Results: Zerio-shot-classification
Table VII shows the experimental results. We compare results based on a strict zero-shot classification setting as in [7]; we only include experimental results that do not use data from downstream tasks during training in our comparisons and gray out all other results as references.
Results show that M2D2 variants achieve top-level performance on all tasks. Among the variants, M2D2- outperforms previous models with a SOTA result of 79.31 for GTZAN, while the second-stage pre-training degrades performance, such as 72.41 with M2D2. M2D2+ brings the ESC-50 and US8K results to the top level, indicating that CLAP training with the limited combination of WavCaps, AudioCaps, and Clotho is advantageous for these tasks compared to M2D2 and M2D2- that use large-scale datasets. Furthermore, M2D2- significantly improves performance on these tasks from M2D20-, indicating the benefit of using an LLM-based text encoder. On the other hand, M2D2 improves CREMA-D closer to the top, CLAP2023.
In summary, we validated that M2D2 achieves top-level performance in the zero-shot classification setting.
@@
| (Proposed models) | |||||||
| M2D2- | 23.15 | 43.43 | 90.07 | 76.52 | 17.41 | 79.31 | 31.76 |
| M2D2 | 28.27 | 46.76 | 93.67 | 81.92 | 23.50 | 72.41 | 24.57 |
| M2D2+ | 27.13 | 42.03 | 94.55 | 82.90 | 21.73 | 70.86 | 29.52 |
| M2D2 | 30.05 | 50.04 | 94.30 | 82.30 | 28.56 | 70.69 | 30.57 |
| (Ablation using a non-LLM text encoder and AudioSet only as audio data) | |||||||
| M2D20-(M2D-CLAP) | 27.24 | 40.82 | 75.45 | 72.40 | 17.73 | 75.17 | 23.39 |
@@
| (Proposed models) | ||||||||||||||||||
| M2D2 | LS+WC | 27.34 | 58.14 | 71.45 | 40.14 | 36.05 | 65.31 | 77.43 | 23.95 | 15.92 | 37.88 | 49.26 | 25.37 | 18.09 | 39.90 | 52.54 | 12.39 | |
| M2D2 | LS+WC+A+C | 37.24 | 72.52 | 84.54 | 52.00 | 52.56 | 82.34 | 90.60 | 38.74 | 18.09 | 42.53 | 55.52 | 28.51 | 24.88 | 48.13 | 60.48 | 15.53 | |
| M2D2+ | WC+A+C | 40.54 | 74.52 | 86.92 | 55.09 | 57.99 | 83.91 | 91.43 | 41.72 | 19.43 | 45.84 | 58.64 | 30.55 | 26.32 | 51.20 | 65.07 | 17.65 | |
| M2D2 | WC+A+C | ✓ | 41.96 | 77.07 | 88.57 | 56.74 | 59.25 | 84.74 | 92.79 | 43.24 | 20.08 | 45.99 | 59.48 | 31.21 | 24.98 | 51.67 | 64.59 | 17.35 |
| (Ablations using a non-LLM text encoder. For the audio data, M2D20-(M2D-CLAP) used AudioSet only while +WC used AudioSet and WavCaps) | ||||||||||||||||||
| M2D20-(M2D-CLAP) | LS | 24.97 | 58.58 | 72.50 | 39.00 | 25.50 | 58.93 | 76.07 | 18.59 | 10.24 | 26.49 | 37.07 | 17.33 | 11.00 | 25.93 | 35.31 | 7.20 | |

### v2

@@ IV-A1 Basic setup
We followed M2D for the basic setup and implemented M2D-CLAP on top of the M2D codebase as a variant of M2D-X[23]. Specifically, the audio encoder is ViT-base[39] with a 768-d feature, input audio duration of 6 s, and patch size of . The audio projector is a single-block transformer encoder with a 768-d feature, a single head, a 768-d feed-forward network latent feature, and a class token. It inputs audio features encoded by the audio encoder and outputs the class token feature as a summarized CLAP audio feature.
The audio samples were preprocessed to a log-scaled mel spectrogram with a sampling frequency of 16,000 Hz, window size of 25 ms, hop size of 10 ms, and 80 mel-spaced frequency bins in the range of 50 to 8000 Hz. We standardized spectrograms with the statistics of the first stage dataset (e.g., average of and standard deviation of ).
@@ IV-A2 Pre-training setup
In the additional stage 1.1, we fine-tuned the first-stage audio encoder in the same way as in the fine-tuning experiments in Section IV-B3. In the additional stage 2.1, we used the same setting as in the second stage once again, except for a masking ratio of 0.0 (i.e., no masking) and a limited dataset.
@@ IV-A3 Pre-training datasets
In the first stage, the pre-training data consists of AudioSet[17], VGGSound[45], and WavCaps[7], 2,532,215 samples in total. We made it closer to CLAP[6]; they used 4.6 M samples from AudioSet, WavCaps, and many other datasets. We used Auto-ACD[46] for the caption data of AudioSet and VGGSound and supplemented the missing captions with ”The sound of labels” using their labels. Since the curated dataset included data samples for the downstream tasks (AudioCaps[41], Clotho[47], ESC-50[16], US8K[48]) used in the evaluations, we removed all of these data samples. For training audio, we randomly cropped 6-s segments.
@@
| (Previous studies: Audio-language models) | |||||||||||
@@
| (Proposed: Audio-language model) | |||||||||||
| M2D-CLAP | ✓ | 97.9 0.5 | 89.7 0.3 | 94.8 0.2 | 56.3 0.5 | 97.1 0.0 | 69.7 0.3 | 86.3 1.8 | 76.7 0.4 | 41.1 0.3 | 78.8 |
| (Ablations: No fine-tuning to AudioSet. M2D-CLAP is pre-trained using a non-LLM text encoder and audio data only from AudioSet.) | |||||||||||
| M2D-CLAP | 95.2 0.3 | 89.5 0.2 | 95.6 0.1 | 68.6 0.5 | 98.1 0.1 | 73.4 0.7 | 86.6 1.6 | 76.5 0.4 | 42.2 0.3 | 80.6 | |
| M2D-CLAP[11] | 96.3 0.3 | 88.8 0.6 | 95.8 0.3 | 70.3 0.4 | 98.3 0.1 | 73.4 0.2 | 84.1 1.5 | 78.0 0.5 | 42.4 0.6 | 80.8 | |
@@ IV-A4 Evaluation tasks
All tasks are classification setups, and results are accuracies or mean average precision (mAP). When data contained variable-length audio, we randomly cropped or zero-padded the average audio length for each sample to create a batch of length-aligned audio fed into the model as in [15][23].
We also used AudioCaps[41] and Clotho[47] for audio-text retrieval and audio captioning. In addition, we used various music tasks supported by the MARBLE benchmark[60].
@@
| (Previous studies: General audio and audio-language models) | |||||||||||||||||
@@
| (Proposed: General audio-language model) | |||||||||||||||||
| M2D-CLAP | 91.8 | 41.6 | 85.9 | 58.2 | 76.2 | 79.7 | 89.5 | 78.9 | 92.7 | 79.7 | 22.2 | 78.8 | 15.8 | 88.3 | 20.9 | 84.0 | 30.9 |
| (Ablations: No fine-tuning to AudioSet. M2D-CLAP is pre-trained using a non-LLM text encoder and audio data only from AudioSet.) | |||||||||||||||||
| M2D-CLAP | 91.5 | 40.8 | 82.6 | 59.3 | 76.7 | 77.4 | 89.1 | 77.4 | 91.8 | 79.3 | 21.7 | 78.0 | 15.9 | 87.9 | 20.8 | 83.7 | 30.3 |
| M2D-CLAP[11] | 91.6 | 41.0 | 84.3 | 61.9 | 77.4 | 80.6 | 88.8 | 78.1 | 92.3 | 79.0 | 21.6 | 78.2 | 16.0 | 88.0 | 20.8 | 83.8 | 30.4 |
IV-B Evaluation: Audio Features
We evaluated the effectiveness of the pre-trained audio features under the frozen feature and fine-tuning settings on general tasks and the frozen feature setting on music tasks. We followed [15][23] for the frozen feature (linear evaluation) and fine-tuning experimental setup using a unified evaluation platform, EVAR444https://github.com/nttcslab/eval-audio-repr, for general tasks. We also evaluated features for music tasks using the MARBLE benchmark[60].
IV-B1 Frozen audio features on general audio tasks
Experimental setup
Results: Frozen audio features on general audio tasks
The linear evaluation results in Table II show that M2D and M2D-CLAP variants perform well across the tasks, whereas supervised learning models (AST, HTS-AT, and PANNs) and CLAP models (e.g., LAION-CLAP) perform poorly on speech tasks and Surge. For instance, their scores range from 7.8 to 18.1 on VC1 for supervised models and 8.4 to 27.2 on Surge for CLAP models. In contrast, M2D and M2D-CLAP variants achieve significantly higher scores, ranging from 52.7 to 73.4 on VC1 and from 40.5 to 42.4 on Surge. Note that the CLAP models employ backbone audio encoders trained with supervised learning, which limits the generalizability of their audio features to tasks outside the scope of their training labels and captions. For example, a label such as ”Male speech, man speaking” does not capture speaker identity, as required in VC1 (speaker identification task).
BEATs and CED, pre-trained with distillation of the previous training cycle models, show a similar trend to supervised learning models. CED shows 35.2 on VC1, and BEATs shows 35.7 on Surge, for example. These models have distilled the pre-trained models on AudioSet, making them focused more on AudioSet data distribution. This specialization may hinder their generalization to tasks involving different data distributions.
In contrast, M2D, M2D-CLAP, and an SSL model, ATST, perform well in all tasks, and M2D-CLAP achieves SOTA performance on ESC-50 at 97.9 and US8K at 89.7. However, fine-tuning on AudioSet may have reduced the sensitivity of the representations to fine-grained speaker differences, causing a performance drop in M2D-CLAP on VC1 (56.3 compared to 73.4 with M2D), although 56.3 still remains a competitive performance. The first-stage model, M2D-CLAP, also shows slightly lower performance on speech tasks and reduces the overall average performance compared to M2D, suggesting that the use of AudioSet captions or label supervision may be the cause.
In addition, M2D-CLAP, which is pre-trained only on AudioSet using GTE-Base[64] (non-LLM-based sentence embedding model) as a text encoder, performs better than M2D-CLAP in most tasks, indicating that the pre-training setting, especially relying solely on AudioSet is beneficial for performing well with frozen audio features. However, as shown in Section IV-C, its CLAP features severely underperform M2D-CLAP.
In summary, M2D-CLAP maintains strong performance as a frozen audio feature extractor on general audio tasks, albeit with compromised performance on speech tasks.
@@ IV-B2 Frozen audio features on music tasks
To further assess the frozen feature, we evaluated models on the music audio tasks with various types of audio signals.
@@
We used the MARBLE benchmark[60], which covers diverse music audio tasks for music tracks, instrument sounds, and vocals. We selected the tasks for which we could reproduce the results with the released code555https://github.com/a43992899/MARBLE-Benchmark. The tasks include music tagging and genre classification on MagnaTagATune (MTT)[65], GTZAN[53], and MTG-Jamendo (MTG)[66], emotion detection on Emomusic[67] and MTG, pitch and instrument (inst) classification on NSynth[54] and MTG, and vocal technique detection (tech) and singer identification on VocalSet[68]. The evaluation metrics are the area under the ROC curve (ROC) and the average precision (AP), except for Emomusic, which uses the determination coefficient between the model’s regression results and human annotations of valence (R2V) and arousal (R2A)[67]. The results are the averages of five tests with different random seeds.
@@
The results of the music task benchmark in Table III show that general audio models perform comparably to the SOTA music audio models. They even outperform music models on GTZAN and NSynth (Inst) while underperforming them in pitch classification NSynth (Pitch). While music models are pre-trained on large-scale music databases, general models are also pre-trained using large-scale music data because more than half (1M) of the AudioSet samples are tagged as ”music”[17]. In addition, most of the music and general models share a similar SSL-based training scheme. These facts would explain the comparable performance of general models on the music-specific benchmark. However, CLAP2023, which uses a supervised learning-based audio backbone, underperforms other general and music models (except NSynth (Inst)666The CLAP2023 training data contains NSynth.), suggesting that SSL on AudioSet is important for generalizing to music tasks, as demonstrated by other SSL-based general audio models.
The models pre-trained using both SSL and supervised learning of AudioSet (M2D-CLAP and BEATs) show stronger performance, and M2D-CLAP demonstrates new SOTA results on MTT, VocalSet, and MTG. However, these models underperform music audio models on the NSynth (Pitch) task, indicating that music models have higher resolution in capturing frequency-related information than general models. Notably, M2D-CLAP achieves performance on par with MULE, a music model, on this task. In addition, as in the evaluation on general tasks in Table II, M2D-CLAP performs better than M2D-CLAP in most tasks, also indicating that pre-training solely on AudioSet is beneficial for performing well with frozen features.
To summarize, the experiment confirmed that M2D-CLAP performs well and shows SOTA performance in music tasks.
@@ IV-B3 Fine-tuning
For fine-tuning, we focus primarily on evaluating M2D-CLAP. We discuss the final M2D-CLAP as an ablation model because it has already been fine-tuned to AudioSet once in the additional stage 1.1.
@@
We followed M2D[23], ATST[32], and Audio-MAE[31] for the fine-tuning procedure. Specifically, a linear classifier was added on top of the pre-trained model to train the entire network. We used SpecAugment[69], Mixup[70, 15], and Random Resize Crop (RRC)[15] for data augmentation, along with Structured Patchout[71] that randomly masks patches during training. We used SGD, LARS, AdamW optimizers, and cosine annealing[72] learning rate scheduling. We used the hyperparameters for each task as listed in Table IV. For Structured Patchout, we used 768-d features calculated by averaging all ViT outputs because Eq. (2) is not applicable to the masked patches. We used the tasks commonly used in previous studies, namely ESC-50, SPCV2, and VC1, the same as in the linear evaluation, plus AS2M and AS20K.
Unlike the linear evaluation covering most models, we limited fine-tuning evaluation to our models, as reproducing previous best results often involves non-trivial tuning.
@@
| (Proposed: Audio-Language model) | |||||
| M2D-CLAP | 49.0 0.1 | 42.1 0.0 | 97.8 0.1 | 98.4 0.0 | 95.5 0.3 |
| (Ablations: M2D-CLAP have been fine-tuned to AudioSet in stage 1.1.) | |||||
| M2D-CLAP | - | 46.7 0.2 | 98.5 0.0 | 98.4 0.1 | 94.0 0.2 |
| M2D-CLAP[11] | 48.5 0.1 | 41.8 0.2 | 97.4 0.2 | 98.3 0.1 | 95.5 0.2 |
@@
Table V shows that M2D-CLAP improves upon the baseline M2D, achieving SOTA performance on AS2M (49.0) and strong results on AS20K (42.1) and ESC-50 (97.8), outperforming the previous top model, BEATsiter3+, except on ESC-50. However, VC1 performance degrades to 95.5, lower than the original M2D’s performance (96.3), suggesting that learning from the CLAP supervision sacrifices fine-grained feature representation.
Compared to ablation M2D-CLAP, M2D-CLAP shows superior performance on all tasks, suggesting that the effect of the semantic supervision of LLM-based sentence embeddings is effective in fine-tuning, but does not contribute as a frozen feature, as confirmed in Section IV-B1.
The final M2D-CLAP, which we handle as ablation, shows improved performance of 46.7 for AS20K and 98.5 for ESC-50, while 94.0 for VC1, a -1.5 pp degradation compared to M2D-CLAP. The degradation of VC1 reaffirms the trade-off caused by fine-tuning under AudioSet’s supervision.
In summary, M2D-CLAP provides top-performing audio features for fine-tuning. Notably, this is achieved by a single pre-training process, unlike other SOTA models that require multiple stages or models. The first stage model M2D-CLAP also shows consistent performance gains across other experiments in Section IV-B.
@@ IV-C1 Zero-shot Classification
We conducted experiments to evaluate the zero-shot classification performance of CLAP features. In this setting, the model, without any task-specific training, predicts the class label of an audio clip by selecting the one whose CLAP text feature has the highest similarity to the CLAP audio feature.
@@
We chose the tasks found in previous papers, which include ESC-50 (ESC), US8K, CREMA-D (CRD), GTZAN (GTZ), NSynth (NS), AudioSet (AS), and FSD50K (FSD). We used AudioSet and FSD50K as references because M2D-CLAP was pre-trained on AudioSet and partially on FSD50K in WavCaps. We conducted experiments using a standard procedure to infer a class label of audio with the closest cosine distance between the features of each test sample and the label’s caption, and we obtained accuracies from the inference results. Table VI summarizes the rules for converting task labels to the captions. We used the evaluation platform EVAR, as in Section IV-B.
Results: Zero-shot classification
Table VII shows the experimental results. We compare results based on a strict zero-shot classification setting as in [7]; we only include experimental results that do not use data from downstream tasks during training in our comparisons, and all other results are grayed out as references.
Results show that M2D-CLAP variants perform well on all tasks, and M2D-CLAP achieves top-level performance in all tasks. Among the ablation variants, M2D-CLAP outperforms previous models with a SOTA result of 79.31 for GTZAN, while the second-stage pre-training degrades performance, such as 72.41 with M2D-CLAP.
M2D-CLAP brings the ESC-50 and US8K results to the top level, indicating that CLAP training with the limited combination of WavCaps, AudioCaps, and Clotho is advantageous for these tasks compared to the models M2D-CLAP that use large-scale datasets. Furthermore, M2D-CLAP significantly improves performance on these tasks from M2D-CLAP[11], indicating the benefit of using an LLM-based text encoder as well as incorporating additional data from VGGSound and WavCaps.
In summary, we validated that M2D-CLAP achieves top-level performance in the zero-shot classification setting.
@@
| (Proposed model) | |||||||
| M2D-CLAP | 30.05 | 50.04 | 94.30 | 82.30 | 28.56 | 70.69 | 30.57 |
| M2D-CLAP | 23.15 | 43.43 | 90.07 | 76.52 | 17.41 | 79.31 | 31.76 |
| M2D-CLAP | 28.27 | 46.76 | 93.67 | 81.92 | 23.50 | 72.41 | 24.57 |
| M2D-CLAP | 27.13 | 42.03 | 94.55 | 82.90 | 21.73 | 70.86 | 29.52 |
| M2D-CLAP[11] | 27.24 | 40.82 | 75.45 | 72.40 | 17.73 | 75.17 | 23.39 |
@@
| GLAP[78] | LS+WC+A+C | 41.7 | - | 86.1 | - | 54.4 | - | 91.1 | - | 19.4 | - | 58.3 | - | 21.8 | - | 61.5 | - | |
| (Proposed model) | ||||||||||||||||||
| M2D-CLAP | WC+A+C | ✓ | 41.96 | 77.07 | 88.57 | 56.74 | 59.25 | 84.74 | 92.79 | 43.24 | 20.08 | 45.99 | 59.48 | 31.21 | 24.98 | 51.67 | 64.59 | 17.35 |
| M2D-CLAP | LS+WC | 27.34 | 58.14 | 71.45 | 40.14 | 36.05 | 65.31 | 77.43 | 23.95 | 15.92 | 37.88 | 49.26 | 25.37 | 18.09 | 39.90 | 52.54 | 12.39 | |
| M2D-CLAP | LS+WC+A+C | 37.24 | 72.52 | 84.54 | 52.00 | 52.56 | 82.34 | 90.60 | 38.74 | 18.09 | 42.53 | 55.52 | 28.51 | 24.88 | 48.13 | 60.48 | 15.53 | |
| M2D-CLAP | WC+A+C | 40.54 | 74.52 | 86.92 | 55.09 | 57.99 | 83.91 | 91.43 | 41.72 | 19.43 | 45.84 | 58.64 | 30.55 | 26.32 | 51.20 | 65.07 | 17.65 | |
| (Ablations using a non-LLM text encoder. For the training data, M2D-CLAP[11] used AudioSet only while +WC used AudioSet and WavCaps.) | ||||||||||||||||||
| M2D-CLAP[11] | LS | 24.97 | 58.58 | 72.50 | 39.00 | 25.50 | 58.93 | 76.07 | 18.59 | 10.24 | 26.49 | 37.07 | 17.33 | 11.00 | 25.93 | 35.31 | 7.20 | |